tripling the greatest of two consecutive even intergers gives the same result as subracting 10 from the lesser even integer. What is the integers?
Answer by kapilsinghi(68)

tripling the greatest of two consecutive even intergers gives the same result as subracting 10 from the lesser even integer. What is the integers?
let the one even number be 2x
so other number lesser than this is 2x - 2

Tripling the greatest i.e (2x)/3
subracting 10 from the lesser even integer 2x - 2 -10
2x - 2 -10 = (2x)/3
2x - 12 = 2x/3
6x - 36 = 2x
4x = 36
x = 9
so the 2 numbers are 18 and 16
kapilsinghi123@gmail.com